OpenStudy (jmansedg):

Write both equations of the following system in the slope intercept form in order to decide how many solutions the system will have. Select the letter that corresponds with your result. 2x + 3y = 12 4x + 6y = 15 A. one solution B. two solutions C. three solutions D. no solution

OpenStudy (jmansedg):

is it b

OpenStudy (vladthepro):

Yep U got it. I wanted to explain the process but I see u got it.

OpenStudy (jmansedg):

what about this one Solve the following system of equations. 2x = 6 3x – y = 9 A. (0, 3) B. (-3, 0) C. (0, 0) D. (3, 0)

OpenStudy (vladthepro):

Do u have a guess?

OpenStudy (jmansedg):

d

OpenStudy (vladthepro):

Yep :)

OpenStudy (jmansedg):

this one Jamie and Nicole are buying streamers for their mother’s surprise birthday party. They need to buy purple and gold streamers. The purple streamers cost $2 per roll and the gold streamers cost $3 per roll. They have $25 to spend on streamers. Which of the following is a possible combination of streamers? A. 1 roll of purple streamer and 8 rolls of gold streamer B. 3 rolls of purple streamer and 6 rolls of gold streamer C. 7 rolls of purple streamer and 4 rolls of gold streamer D. 9 rolls of purple streamer and 3 rolls of gold streamer

OpenStudy (vladthepro):

3*2 + 6*3 = 24 which is less than $25 So I say B.

OpenStudy (jmansedg):

i was going to say the same thing for my guess

OpenStudy (jmansedg):

Solve the following system of equations. x + y = 5 y = 3 A. (-2, 3) B. (2, -3) C. (0, 0) D. (2, 3)

OpenStudy (jmansedg):

i think d

OpenStudy (vladthepro):

yep

OpenStudy (jmansedg):

Andrew wants to buy Stephanie balloons and roses for Valentine’s Day. Roses cost $2.50 each and balloons are $1.25 each. He has $15 to spend. Which of the following is a possible combination of roses and balloons? A. 1 rose and 11 balloons B. 3 roses and 7 balloons C. 5 roses and 2 balloons D. 6 rose and 1 balloon @VladthePro

OpenStudy (jmansedg):

i think d

OpenStudy (jmansedg):

no c

OpenStudy (jmansedg):

i got an 80% on it so thanks man

OpenStudy (vladthepro):

Wait what abt this lastone?

OpenStudy (jmansedg):

i gussed c and it right

OpenStudy (jmansedg):

is

OpenStudy (vladthepro):

K :) Good job bro

OpenStudy (jmansedg):

thanks you helped me so it was you man